# A Longitudinal Study of Child Wellbeing Assessment via Online Interactions with a Social Robot

### v2

@@ Abstract
Socially Assistive Robots are studied in different Child-Robot Interaction settings. However, logistical constraints limit accessibility, particularly affecting timely support for mental wellbeing. In this work, we have investigated whether online interactions with a robot can be used for the assessment of mental wellbeing in children. The children (N=40, 20 girls and 20 boys; 8-13 years) interacted with the Nao robot (30-45 mins) over three sessions, at least a week apart. Audio-visual recordings were collected throughout the sessions that concluded with the children answering user perception questionnaires pertaining to their anxiety towards the robot, and the robot’s abilities. We divided the participants into three wellbeing clusters (low, med and high tertiles) using their responses to the Short Moods and Feelings Questionnaire (SMFQ) and further analysed how their wellbeing and their perceptions of the robot changed over the wellbeing tertiles, across sessions and across participants’ gender. Our primary findings suggest that (I) online mediated-interactions with robots can be effective in assessing children’s mental wellbeing over time, and (II) children’s overall perception of the robot either improved or remained consistent across time. Supplementary exploratory analyses have also revealed that gender affected the children’s wellbeing assessments as well as their perceptions of the robot.
@@
Keywords Child–robot Interaction  $\cdot$ Wellbeing  $\cdot$ Longitudinal  $\cdot$ Human–robot Interaction  $\cdot$ Psychological Assessment
1 Introduction
@@
Research into Socially Assistive Robots (SARs) in children’s lives have become increasingly prevalent, addressing their potential roles in education [1, 2], autism interventions [3] and more recently in the field of mental health [4, 5]. Children have been shown to consider robots as companions and peers because of the non-intimidating, engaging nature of SARs that provide an interactive, responsive and non-judgemental platform for children to feel comfortable [6, 7]. However, interacting with SARs usually involves visiting a laboratory and thus providing logistical restrictions in availability and accessibility. In sensitive domains like mental health, emotional stigma, shame and attitudes towards seeking help (for example, from therapists [8], online support initiatives [9] and wellbeing applications [10]) may affect self-disclosure in children [11, 12].In addition, children’s inclinations to openly express their emotions are also heavily influenced by their gender as boys have been observed to show increased hesitation towards disclosing their genuine feelings compared to girls [13].
SARs have been shown to help with mood estimation [14] and improvement [15], alleviating anxiety during painful medical procedures [16] and identifying bullying instances among children [17]. However, most of these studies have involved visits to unfamiliar environments like university laboratories or hospitals that might affect children’s interactions and comfort levels. Studies that have investigated the interaction of children with SARs in familiar settings have primarily been in the field of edutainment [1] or autism research [3]. Thus, having a robotic platform that children can interact with in an environment familiar to them, and that can take into account any predispositions in their attitudes, especially in the context of mental wellbeing assessment, might help them to be more candid about their feelings and experiences.
@@
In the context of long-term HRI, SARs have been shown to provide emotional support and companionship to the elderly who formed strong bonds with the robot over long-term interactions [18]. Even online longitudinal interaction with robots has been shown to enhance mood, mitigate feelings of loneliness and promote self-disclosure in adults [15]. Thus, in this work, we present the first study that has explored whether longitudinal online interactions with a robot can be used for the assessment of mental wellbeing in children. Further, we have explored whether co-location affects robotised wellbeing assessment by comparing our research findings with our prior investigation [4] that has involved SARs for the assessment of mental wellbeing through an in-lab setting. We also provide an exploratory supplementary investigation into whether the gender of the participants affects the longitudinal assessment of mental wellbeing of children. For this purpose, we have gathered data from 40 children (20 girls and 20 boys, 8-13 years old) who interacted with a Nao robot (SoftBank Robotics), over the course of three sessions for about 30-45 mins in each session. Our primary findings suggest that (i) online interactions with the robot (using Zoom or Skype) can be effective in the assessment of mental wellbeing of children, and (ii) user perception of the robot largely improved or remained unchanged throughout the course of the three online sessions and was affected by the physical presence of the robot. Our supplementary exploratory analyses have revealed that gender-based variability has been observed in terms of the longitudinal robot-assisted assessments as well as the children’s perception of robot’s behaviour and capabilities. We found that robot-assisted assessment was able to amplify the differences across varying levels of wellbeing for both boys and girls for the first session and only for boys for the second session. We have also observed that girls had a higher opinion of the robot as their confidante as compared to boys. We hope that the findings from this work can provide a valuable stepping-stone for upcoming robot-led initiatives that are more accessible to children for the assessment of their mental wellbeing.
2 Related Works
2.1 Wellbeing Assessments
@@
However, these initiatives rely heavily on the assumption that the recorded responses of the children are authentic and representative of their true feelings [24]. Children’s responses might be influenced by their reluctance to divulge sensitive information about themselves or being intimidated by the presence of other adults in the room whom they might consider authority figures [25]. As such, they may provide socially desirable answers that might not reflect their genuine thoughts and emotions [24, 26, 27]. Previous studies have shown that people gradually open up to robots over time [15], with those experiencing negative emotional states being more likely to share extensively with robots [28]. Furthermore, SARs have been shown to be very effective in establishing trust particularly with children [29], who have often attributed robots as peers and friends [30, 31]. As such, robots could potentially be a valuable addition to children’s mental wellbeing assessment pipeline (Figure 1 (b)) possibly eliciting more candid responses from the children that accurately represent their feelings. In addition, robots’ multimodal sensing capabilities can be leveraged to identify non-verbal cues that extends children’s verbal responses, garnering insights from children’s behaviour such as their facial expressions and vocal tone. Consequently, critical observations can be flagged for child mental health practitioners to facilitate a diagnosis so that necessary support, if needed, can be provided at an early stage [see similar discussion in 32].
@@
Figure 1: (a) Pipeline showing the procedure for mental wellbeing assessment of children in the UK [21, 22, 23]. This pipeline is representative of certain assessment scenarios and is not standardised across all clinical and research assessments conducted in the UK. (b) Proposed robotised mental health assessment pipeline.
2.2 Child robot interaction (CRI) in mental health.
@@
Exploring health and wellbeing applications has been an emerging frontier in CRI literature. Studies have focussed on how SARs have been used to provide emotional support for a variety of health and wellness-related situations [14, 16, 33, 34, 17, 4]. For example, Gamborino et al [14] have used SARs for the assessment and enhancement of mood in children. Rossi et al [16] have investigated how robots have been used to alleviate stress and anxiety in children in hospital consultations. Henkemans et al [33] have used robots to provide motivation to children for self-management of diabetes. Meghdari et al [34] have observed that social robots provide comfort to pediatric cancer patients. Robots have also been used to investigate cases of bullying in children [17]. More recently, SARs have been successfully employed for the evaluation of mental wellbeing in children [4]. However, all the aforementioned works have involved children’s physical presence in unfamiliar settings like hospitals and research laboratories which might affect children’s comfort levels and also contribute towards their hesitancy in opening up.
2.3 Longitudinal CRI.
Long-term interactions facilitate an evaluation of the influence of novelty effect [35] on the child-robot interaction experience. For instance, Kozima et al [36] have observed that long-term interactions with robots have elicited pro-social behaviour among preschoolers. Abdelmohsen et al [37] have shown that longitudinal sessions with virtual robots have improved conversational skills in children with autism. Russel et al [38] have shown how robot-assisted intervention has helped in the reduction of stress and better coping among children with cystic fibrosis. Studies have also shown how the attention and engagement of children decrease over time upon prolonged interaction with robots. For example, Serholt et al [39] have shown that the social responses from children had reduced over multiple interactions with their robot tutors. Salter et al [40] reported that children showed increased boredom after prolonged exposure to the robots in their schools. However, none of these works have investigated whether robots can be used for longitudinal assessment of children’s mental wellbeing.
2.4 Remote access to robots.
@@
Having remote access to SARs might help in reducing the stress among participants that can be attributed to unfamiliar environments such as research labs and hospitals. For example, in adults, Laban et al [15] have observed that interactions with online mediated robots have promoted enhancement of mood and self-disclosure among their participants and that adults who are experiencing negative emotions tend to disclose more to a social robot despite being remote [28]. In the case of children, speaking with a social robot online has encouraged diabetic children to maintain a diary [41]. Lytridis et al [42] have used robots to provide personalised support and special education to children with ASD, without in-person interaction with the robot. Pasupuleti et al [43] have employed teleconferencing with a robot to promote good hand hygiene among children. Calvo et al [44] have demonstrated how repeated exposure to online social robots promotes trust among children in a storytelling task. Previous HRI studies have also reported that co-location with the robot does not necessarily improve the objective of the experimental tasks under consideration. For example, Gittens et al [45] have shown that conducting HRI sessions online does not negatively affect the interaction experience. Kennedy et al [46] have shown that the robot’s physical presence did not improve children’s learning. Hence, having a remote robot platform can mitigate the logistical constraints associated with CRI, enabling better accessibility, and thus, encouraging timely support in sensitive matters related to children’s mental wellbeing.
2.5 Gender differences in mental health and HRI.
@@
With regard to mental health, gender-based variability has been observed in children about their perceived wellbeing. For example, societal constructs concerning masculinity have been shown to dictate boys’ attitudes towards mental health and seeking help when in need [47]. Boys have also shown to have decreased awareness in relation to their mental health [13] and have increased stigma associated with mental health concerns like depression as compared to girls [48]. Even in the field of HRI, the gender of the participants seemed to influence their perceived perception of the robot and their interaction experience. For example, Mutlu et al [49] have shown that the male participants have reported higher self-perceived positive affect when interacting with a robot as compared to the female participants. Further, Strait et al [50] showed a difference in perceived ratings between males and females in relation to changes in verbal and non-verbal behaviours of the robot (i.e. women reacted favourably to increased politeness from the robot). Therefore, in order to develop a holistic understanding of the mental wellbeing of the current and future generations, it is important to account for gender-related differences when designing robot-assisted assessment tools for children.
2.6 Motivation for using online mediated robots
In the study described in this paper, we have used online mediated robots, wherein the participants communicated with the physical robot via computer mediated communication (CMC) such as Zoom or Skype. As discussions about the potential applications of social robots in well-being settings gain prominence, it is crucial to acknowledge that physical HRI research often relies on laboratory-constrained environments. This reliance limits the generalization of results, interactions, and insights across diverse populations and conditions. The COVID-19 pandemic, by accelerating the adoption of computer-mediated communication in households [51], has opened new avenues for HRI research. This shift towards virtual and CMC methods enables researchers to explore the interaction with robots in home environments via personal screens [52], allowing for rigorous and reproducible research that effectively supports individuals’ well-being in their natural settings. For instance, CMC interactions with social robots have been leveraged to support informal caregivers [53], who often struggle to find time and energy for self-care, leading to overlooked health issues [54]. This approach offers them the opportunity to discuss their concerns with an online-mediated robot, ensuring continuous support in their natural environments. These strategies not only enhance the accessibility of emotional well-being interventions to a wider range of populations, thereby promoting inclusivity and diversity but also enable the generalization of HRI research findings to diverse groups that may be restricted from participating in in-person studies. Despite concerns about the loss of physical embodiment, recent studies have found no significant differences in participants’ perceptions and interactions with robots in virtual versus in-person settings [55, 56, 57, 58], providing strong evidence for the efficacy of CMC as an alternative means for HRI experimentation, especially in unique settings such as CRI with children from around the world.
Our rationale for using this approach over other remote methods (such as chatbots, in-home robots or virtual animated characters of robots) that also aim to provide increased accessibility to the robots, was driven by the following factors. Firstly, while chatbots have become increasingly popular in the mental health domain, often being used in mental health applications, they require some initiative from the participants [59]. This would be difficult to encourage in our study as the participants were children and could even lead to increased drop-out rates in the longitudinal study conducted here. Secondly, in the case of in-home robots, there is a logistical disadvantage in terms of accessibility to every child  [^2] https://automationswitch.com/disadvantages-of-domestic-robots/, present in different parts of the world in comparison with smart devices required for online mediated robots that are more readily available in households  [^3] https://explodingtopics.com/blog/smartphone-stats. As such, there would be monetary costs associated with purchasing either on the researchers’ end or on the participants’, creating inconvenience for each party. Thirdly, prior research [60] has indicated that physical presence and not physical embodiment of the robots affect user responses. In other words, Li et al [60] have demonstrated that co-present robots (present in the same room as the participants) have been shown to affect users’ behaviour in comparison with virtual (animated character of robots) or telepresent/online (physical robots present on a camera feed). However, there was no significant effect between user responses to virtual robots over telepresent robots, thus, virtual animated robots do not offer any notable advantage. Fourthly, research has also shown that self-disclosure and mood were improved when participants interacted with the Pepper robot online using Zoom [15, 53, 61]. Fifthly, children have also become very familiar with videoconferencing tools such as Zoom or Skype, during the Covid-19 pandemic, thus, providing ease of convenience and familiarity when interacting with online mediated robots [^4] https://www.zoom.com/en/blog/how-zoom-supported-young-people-in-uk-during-pandemic/. Finally, we also wanted to make valid comparisons with our previous works conducted in [4]. As such, we followed the experimental protocol conducted in [4] and [62] and only modified the mode of interaction from in-person to online. Using a virtual robot (or agent) would have changed the robot’s form and appearance to an animated character making it difficult to ascertain if any changes in findings were due to the mode of the interaction (in-person vs. online) or the robot’s form (virtual animated character or the actual physical robot). Spitale et al [63] have shown that in HRI for wellbeing studies changing robot form leads to significant changes in user perception and interaction. To this end, we have employed robots that interacted with the participants through Zoom or Skype, to explore how they affect the assessment of mental wellbeing in children.
3 Research Questions and Contributions

### v3

@@ Abstract.
Socially Assistive Robots are studied in different Child-Robot Interaction settings. However, logistical constraints limit accessibility, particularly affecting timely support for mental wellbeing. In this work, we have investigated whether online interactions with a robot can be used for the assessment of mental wellbeing in children. The children (N=40, 20 girls and 20 boys; 8-13 years) interacted with the Nao robot (30-45 mins) over three sessions, at least a week apart. Audio-visual recordings were collected throughout the sessions that concluded with the children answering user perception questionnaires pertaining to their anxiety towards the robot, and the robot’s abilities. We divided the participants into three wellbeing clusters (low, med and high tertiles) using their responses to the Short Moods and Feelings Questionnaire (SMFQ) and further analysed how their wellbeing and their perceptions of the robot changed over the wellbeing tertiles, across sessions and across participants’ gender. Our primary findings suggest that (I) online mediated-interactions with robots can be effective in assessing children’s mental wellbeing over time, and (II) children’s overall perception of the robot either improved or remained consistent across time. Supplementary exploratory analyses have also revealed that the gender of the children affected their wellbeing assessments with interactions effectively distinguishing between varying levels of wellbeing for both boys and girls for the first session and only for boys during the second session. The analyses have also revealed that girls have a higher opinion of the robot as a confidante as compared with boys. Findings from this work affirm the potential of using online mediated interactions with robots for the assessment of the mental wellbeing of children.
@@
Robotised assessment, Mental wellbeing, Child-robot interaction, longitudinal, human-robot interaction
1. Introduction
@@
Research into Socially Assistive Robots (SARs) in children’s lives have become increasingly prevalent, addressing their potential roles in education (Tanaka et al., 2015; Henschel et al., 2021), autism interventions (Scassellati et al., 2018) and more recently in the field of mental health (Abbasi et al., 2022, 2024b). Children have been shown to consider robots as companions and peers because of the non-intimidating, engaging nature of SARs that provide an interactive, responsive and non-judgemental platform for children to feel comfortable (Kory-Westlund and Breazeal, 2019; De Haas et al., 2020). However, interacting with SARs usually involves visiting a laboratory and thus providing logistical restrictions in availability and accessibility. In sensitive domains like mental health, emotional stigma, shame and attitudes towards seeking help (for example, from therapists (Thepthien et al., 2019), online support initiatives (Yang et al., 2019) and wellbeing applications (Punukollu and Marques, 2019a)) may affect self-disclosure in children (Smart and Wegner, 2000; Corrigan et al., 2014).In addition, children’s inclinations to openly express their emotions are also heavily influenced by their gender as boys have been observed to show increased hesitation towards disclosing their genuine feelings compared to girls (Chandra and Minkovitz, 2006).
SARs have been shown to help with mood estimation (Gamborino et al., 2019) and improvement (Laban et al., 2024), alleviating anxiety during painful medical procedures (Rossi et al., 2022) and identifying bullying instances among children (Bethel et al., 2016). However, most of these studies have involved visits to unfamiliar environments like university laboratories or hospitals that might affect children’s interactions and comfort levels. Studies that have investigated the interaction of children with SARs in familiar settings have primarily been in the field of edutainment (Tanaka et al., 2015) or autism research (Scassellati et al., 2018). As discussions about the potential applications of social robots in well-being settings gain prominence, it is crucial to acknowledge that physical HRI research often relies on laboratory-constrained environments. This reliance limits the generalization of results, interactions, and insights across diverse populations and conditions. The COVID-19 pandemic, by accelerating the adoption of computer-mediated communication in households (Choi and Choung, 2021), has opened new avenues for HRI research. This shift towards virtual and computer-mediated communication (CMC) methods enables researchers to explore the interaction with robots in home environments via personal screens (Feil-Seifer et al., 2020), allowing for rigorous and reproducible research that effectively supports individuals’ well-being in their natural settings. For instance, CMC interactions with social robots have been leveraged to support informal caregivers (Laban et al., 2025), who often struggle to find time and energy for self-care, leading to overlooked health issues (Revenson et al., 2016). This approach offers them the opportunity to discuss their concerns with an online-mediated robot, ensuring continuous support in their natural environments. These strategies not only enhance the accessibility of emotional well-being interventions to a wider range of populations, thereby promoting inclusivity and diversity but also enable the generalization of HRI research findings to diverse groups that may be restricted from participating in in-person studies. Despite concerns about the loss of physical embodiment, recent studies have found no significant differences in participants’ perceptions and interactions with robots in virtual versus in-person settings (Laban et al., 2021; Honig and Oron-Gilad, 2020; Gittens, 2021; Gittens and Garnes, 2022a), providing strong evidence for the efficacy of CMC as an alternative means for HRI experimentation, especially in unique settings such as CRI with children from around the world. Thus, having a robotic platform that children can interact with in an environment familiar to them, and that can take into account any predispositions in their attitudes, especially in the context of mental wellbeing assessment, might help them to be more candid about their feelings and experiences.
@@
In the context of long-term HRI, SARs have been shown to provide emotional support and companionship to the elderly who formed strong bonds with the robot over long-term interactions (Yamazaki et al., 2023). Even online longitudinal interaction with robots has been shown to enhance mood, mitigate feelings of loneliness and promote self-disclosure in adults (Laban et al., 2024). Thus, in this work, we present the first study that has explored whether longitudinal online interactions with a robot can be used for the assessment of mental wellbeing in children. Further, we have explored whether co-location affects robotised wellbeing assessment by comparing our research findings with our prior investigation (Abbasi et al., 2022, 2024b) that has involved SARs for the assessment of mental wellbeing through an in-lab setting. We also provide an exploratory supplementary investigation into whether the gender of the participants affects the longitudinal assessment of mental wellbeing of children. For this purpose, we have gathered data from 40 children (20 girls and 20 boys, 8-13 years old) who interacted with a Nao robot (SoftBank Robotics), over the course of three sessions for about 30-45 mins in each session. Our primary findings suggest that (i) online interactions with the robot (using Zoom or Skype) can be effective in the assessment of mental wellbeing of children, and (ii) user perception of the robot largely improved or remained unchanged throughout the course of the three online sessions and was affected by the physical presence of the robot especially in the case of anxiety attributed towards the robot and the perceived intelligence of the robot. Our supplementary exploratory analyses have revealed that gender-based variability has been observed in terms of the longitudinal robot-assisted assessments as well as the children’s perception of robot’s behaviour and capabilities. We found that robot-assisted assessment was able to amplify the differences across varying levels of wellbeing for both boys and girls for the first session and only for boys for the second session. We have also observed that girls had a higher opinion of the robot as their confidante as compared to boys. In summary, this work builds upon our prior work where we observed children interacting with a robot in an in-lab setting (Abbasi et al., 2022, 2024b). As mentioned above, our findings indicate that even in an online setting, robotised frameworks can be used to effectively differentiate between children belonging to varying levels of wellbeing. This consistency reinforces the adaptability and robustness of our approach paving the way for broader applications, across different settings, in assessing children’s wellbeing. We hope that the findings from this work can provide a valuable stepping-stone for upcoming robot-led initiatives that are more accessible to children for the assessment of their mental wellbeing.
2. Related Works
2.1. Wellbeing Assessments
@@
However, these initiatives rely heavily on the assumption that the recorded responses of the children are authentic and representative of their true feelings (Tourangeau et al., 2000). Children’s responses might be influenced by their reluctance to divulge sensitive information about themselves or being intimidated by the presence of other adults in the room whom they might consider authority figures (O’Reilly and et al., 2015). As such, they may provide socially desirable answers that might not reflect their genuine thoughts and emotions (Tourangeau et al., 2000; Nock and Kazdin, 2001; Godoi et al., 2020). Previous studies have shown that people gradually open up to robots over time (Laban et al., 2024), with those experiencing negative emotional states being more likely to share extensively with robots (Laban et al., 2023). Furthermore, SARs have been shown to be very effective in establishing trust particularly with children (Di Dio et al., 2020), who have often attributed robots as peers and friends (Fior et al., 2010; Kanda et al., 2004). As such, robots could potentially be a valuable addition to children’s mental wellbeing assessment pipeline (Figure 1 (b)) possibly eliciting more candid responses from the children that accurately represent their feelings. In addition, robots’ multimodal sensing capabilities can be leveraged to identify non-verbal cues that extends children’s verbal responses, garnering insights from children’s behaviour such as their facial expressions and vocal tone. Consequently, critical observations can be flagged for child mental health practitioners to facilitate a diagnosis so that necessary support, if needed, can be provided at an early stage (see similar discussion in Laban et al., 2022a).
@@
Figure 1. (a) Pipeline showing the procedure for mental wellbeing assessment of children in the UK (Ford et al., 2020; Mansfield et al., 2020; Burn et al., 2022). This pipeline is representative of certain assessment scenarios and is not standardised across all clinical and research assessments conducted in the UK. (b) Proposed robotised mental health assessment pipeline.
2.2. Digital Mental Health Support
In recent years, mobile applications have emerged as a cost-effective and accessible means of delivering mental health interventions, offering a wide range of support to users on their personal devices. For instance, apps providing cognitive behavioral therapy (CBT), mood tracking, and emotional regulation strategies have gained popularity for their ability to scale and deliver interventions to underserved populations (Guracho et al., 2023). Mobile apps specifically designed for children, such as those targeting emotional self-awareness and mood monitoring, have shown potential to engage young users and offer additional support outside of traditional clinical settings (Punukollu and Marques, 2019b). However, concerns regarding the lack of regulatory oversight, as well as limited evidence on the effectiveness and safety of many commercially available apps, present significant challenges in their widespread adoption and clinical recommendation (Grist et al., 2017). Socially Assistive Robots (SARs) present an alternative approach, offering a unique advantage through their physical cues of embodiment and interactive capabilities (Laban et al., 2022a). Unlike mobile apps, SARs engage users in multimodal interactions, incorporating voice, gestures, and facial expressions to foster rapport and elicit richer self-disclosures (Laban et al., 2022a; Spitale and Gunes, 2022; Spitale et al., 2024; Laban, 2024). The embodied nature of SARs allows for more naturalistic social interactions, which is particularly beneficial in mental health settings where trust and empathy are critical (Cano et al., 2021). This interactive and dynamic form of engagement is especially important when working with children, who often require a playful and empathetic approach to feel comfortable discussing their mental state (Rudenko et al., 2024). Therefore, the use of SARs in this context can be a powerful tool for clinicians aiming to assess and support the emotional well-being of children in a non-threatening, child-friendly manner.
2.3. Child robot interaction (CRI) in mental health.
@@ 2.2. Digital Mental Health Support
Exploring health and wellbeing applications has been an emerging frontier in CRI literature. Studies have focussed on how SARs have been used to provide emotional support for a variety of health and wellness-related situations (Gamborino et al., 2019; Rossi et al., 2022; Henkemans et al., 2017; Meghdari et al., 2018; Bethel et al., 2016; Abbasi et al., 2022). For example, Gamborino et al (Gamborino et al., 2019) have used SARs for the assessment and enhancement of mood in children. Rossi et al (Rossi et al., 2022) have investigated how robots have been used to alleviate stress and anxiety in children in hospital consultations. Henkemans et al (Henkemans et al., 2017) have used robots to provide motivation to children for self-management of diabetes. Meghdari et al (Meghdari et al., 2018) have observed that social robots provide comfort to pediatric cancer patients. Robots have also been used to investigate cases of bullying in children (Bethel et al., 2016). More recently, SARs have been successfully employed for the evaluation of mental wellbeing in children (Abbasi et al., 2022). However, all the aforementioned works have involved children’s physical presence in unfamiliar settings like hospitals and research laboratories which might affect children’s comfort levels and also contribute towards their hesitancy in opening up.
2.4. Remote access to robots.
@@
Having remote access to SARs might help in reducing the stress among participants that can be attributed to unfamiliar environments such as research labs and hospitals. For example, in adults, Laban et al (Laban et al., 2024) have observed that interactions with online mediated robots have promoted enhancement of mood and self-disclosure among their participants and that adults who are experiencing negative emotions tend to disclose more to a social robot despite being remote (Laban et al., 2023). In the case of children, speaking with a social robot online has encouraged diabetic children to maintain a diary (Van Der Drift et al., 2014). Lytridis et al (Lytridis et al., 2020) have used robots to provide personalised support and special education to children with ASD, without in-person interaction with the robot. Pasupuleti et al (Pasupuleti et al., 2022) have employed teleconferencing with a robot to promote good hand hygiene among children. Calvo et al (Calvo-Barajas and Castellano, 2022) have demonstrated how repeated exposure to online social robots promotes trust among children in a storytelling task. Previous HRI studies have also reported that co-location with the robot does not necessarily improve the objective of the experimental tasks under consideration. For example, Gittens et al (Gittens and Garnes, 2022b) have shown that conducting HRI sessions online does not negatively affect the interaction experience. Kennedy et al (Kennedy et al., 2015) have shown that the robot’s physical presence did not improve children’s learning. Hence, having a remote robot platform can mitigate the logistical constraints associated with CRI, enabling better accessibility for potential longitudinal interactions, and thus encouraging timely and consistent support in sensitive matters related to children’s mental wellbeing.
2.5. Longitudinal CRI.
Long-term interactions facilitate an evaluation of the influence of novelty effect (Smedegaard, 2019) on the child-robot interaction experience. For instance, Kozima et al (Kozima and Nakagawa, 2007) have observed that long-term interactions with robots have elicited pro-social behaviour among preschoolers. Abdelmohsen et al (Abdelmohsen and Arafa, 2021) have shown that longitudinal sessions with virtual robots have improved conversational skills in children with autism. Russel et al (Russell et al., 2021) have shown how robot-assisted intervention has helped in the reduction of stress and better coping among children with cystic fibrosis. Studies have also shown how the attention and engagement of children decrease over time upon prolonged interaction with robots. For example, Serholt et al (Serholt and Barendregt, 2016) have shown that the social responses from children had reduced over multiple interactions with their robot tutors. Salter et al (Salter et al., 2004) reported that children showed increased boredom after prolonged exposure to the robots in their schools. As such, understanding the long-term interactions between robots and children can provide compelling insights into developing personalised intervention strategies that account for individual variability, especially in the context of mental wellbeing assessments.
2.6. Gender differences in mental health and HRI.
@@
With regard to mental health, gender-based variability has been observed in children about their perceived wellbeing. For example, societal constructs concerning masculinity have been shown to dictate boys’ attitudes towards mental health and seeking help when in need (Pearson, 2023). Boys have also shown to have decreased awareness in relation to their mental health (Chandra and Minkovitz, 2006) and have increased stigma associated with mental health concerns like depression as compared to girls (Lindsey et al., 2010). Even in the field of HRI, the gender of the participants seemed to influence their perceived perception of the robot and their interaction experience. For example, Mutlu et al (Mutlu et al., 2006) have shown that the male participants have reported higher self-perceived positive affect when interacting with a robot as compared to the female participants. Further, Strait et al (Strait et al., 2015) showed a difference in perceived ratings between males and females in relation to changes in verbal and non-verbal behaviours of the robot (i.e. women reacted favourably to increased politeness from the robot). Therefore, in order to develop a holistic understanding of the mental wellbeing of the current and future generations, it is important to account for gender-related differences when designing robot-assisted assessment tools for children.
2.7. Research Questions and Contributions